# Nonlinear Propagation of Non-Gaussian Uncertainties: Theory and Applications to Spacecraft Trajectory Design

### v1

@@ 3.1.2 Perturbed Case
$\begin{split}\dot{\boldsymbol{r}}&=\boldsymbol{v}\\ \dot{v}_{x}&=-\frac{\mu}{r^{3}}x-\dfrac{3\mu J_{2}R_{e}^{2}}{2r^{5}}\bigg{(}1-\dfrac{5z^{2}}{r^{2}}\bigg{)}x\\ \dot{v}_{y}&=-\frac{\mu}{r^{3}}y-\dfrac{3\mu J_{2}R_{e}^{2}}{2r^{5}}\bigg{(}1-\dfrac{5z^{2}}{r^{2}}\bigg{)}y\\ \dot{v}_{z}&=-\frac{\mu}{r^{3}}z-\dfrac{3\mu J_{2}R_{e}^{2}}{2r^{5}}\bigg{(}3-\dfrac{5z^{2}}{r^{2}}\bigg{)}z\text{,}\end{split}$ (26)
@@
As for the example in Sec. 3.1.1, also in this case the coupled equations of motion, cause the initial uncertainties in the $x,y$ component and in the parameters to eventually affect all six components of the state during the propagation. We use the Monte Carlo propagation with 10 million samples as ground truth, and we compare the reconstructed statistical moments with those found with the proposed approach, up to the Taylor expansion of order 5. As before, the time required to compute the Taylor expansion at order five of the state is orders of magnitude lower than the cost to perform the Monte Carlo. On the left of Fig. 4, we display how the accuracy of the reconstruction in the first, second, and third moment behaves as a function of time, using the absolute and relative errors in the moment reconstruction discussed in Eq. (24). The error is always maintained below 1% of accuracy, for all the first three moments tensors. Furthermore, as seen from the right figure, order 1, which consists of just using the first-order state transition matrix, does not accurately reconstruct the second moment. Instead, increasing the order to two, three, and four, brings substantial improvements in the error of the reconstruction, while between order four and order five there is not a significant gap.
Figure 4: Left: error (absolute on top and relative below) of reconstruction (between the proposed method and Monte Carlo) in the first three statistical moments as a function of propagation time; center: (absolute on top and relative below) error of reconstruction of the second moment, for different orders, as a function of time.
@@ 3.2.1 Taylor Polynomial Convergence Radius of Different Families
Table 2: Initial conditions, period, mass ratio parameter, and stability indices of seven different periodic orbits used in the experiments, within the Earth-Moon and Saturn-Enceladus CR3BP system.
@@
Figure 5: Three-dimensional CR3BP periodic orbits in the Earth-Moon (upper left) and Enceladus-Saturn (upper center) systems, together with the Cauchy-Hadamard’s radius of convergence estimation, at different orders, for each periodic orbit.
To generate these, for each orbit, we construct the Taylor expansion after a period and compute the Cauchy-Hadamard radius of convergence for each component of the state. Then, the lowest radius across the different components estimated at the highest order can be used as a proxy of the ball radius in which the Taylor series converges. Hence, this guides the choice for the initial uncertainty size that can be acceptable before the Taylor convergence loses its accuracy. While this criterion is often used in practice, it is important to remember that it would only be valid if the Taylor series is expanded to infinite terms: truncating at lower orders only results in an approximation of the convergence radius. To verify that the criterion is working properly, in Fig. 6, we also report two projections of the ground truth Poincaré map, propagated at different ball radii, together with the associated error w.r.t. the Taylor approximation, for all the different periodic orbits (each row corresponds to a periodic orbit). In the figure, we use the same color for ball radii of the same size. As we observe, the error of the Taylor approximation is maintained very low, when the perturbation is below the Cauchy-Hadamard’s radius of convergence. We also perform the exact same analysis for the Saturn-Enceladus short-period L4 orbit, which we report in Fig. 7. Also in this case, we confirm that for perturbations below the Cauchy-Hadamard’s radius of convergence, the errors between the Taylor approximation and the ground truth orbit are maintained very low, thereby confirming that this criterion can be used as a good proxy for the region of validity of the Taylor expansion.
Figure 6: Numerically integrated (i.e., ground truth) Poincaré maps, for different perturbation radii from the nominal trajectory (each color belongs to a perturbation), together with the error between the ground truth and the Taylor approximation.
@@ 3.2.2 Statistical Moments Mapping onto Poincaré maps
Figure 8: Two-dimensional projections of the propagated statistics of the moments at the Poincaré map, for different Taylor approximation orders (from 1 to 5) as compared with the ground truth statistics (computed with Monte Carlo).
@@
Figure 8 compares eight different projections of the state components for the estimated covariance ellipsoid with five different orders of the Taylor series (ranging from order 1 to order 5) against the statistical moments computed with a 10 million-sample Monte Carlo simulation (considered the ground truth). In the same figure, we also display the initial (in blue) and final (in orange) Monte Carlo samples, for the uncertain coordinates. As we observe qualitatively, the accuracy of the reconstruction of the covariance ellipsoid improves as the order is increased. To confirm this, we compute the relative error between the reconstructed and ground truth second order moment using Eq. (24), obtaining for order 1 $\varepsilon_{r,2}=2.4017$, for order 2 $\varepsilon_{r,2}=0.0109$, for order 3 $\varepsilon_{r,2}=0.0106$, for order 4, and order 5, $\varepsilon_{r,2}=0.0105$. This confirms not only that the accuracy improves, but also that for this case there is not an improvement between order 4 and order 5, which suggests that the Taylor approximation could have been expanded only up to order 4, obtaining the same accuracy, if the objective was the reconstruction of the second moment only.
@@ 4 Conclusion
In this paper, we present a novel approach to propagate uncertainties via high-order Taylor expansions of the flow and moment generating functions. Although prior work in the literature focuses on Gaussian distributions, we exploit the relationship between moment generating functions and distribution moments to extend high-order uncertainty propagation techniques to non-Gaussian scenarios, thereby significantly broadening their applicability across a wide range of problems and uncertainty types.
The proposed approach is general by design and applicable to arbitrary dynamical systems as shown by numerical simulations that cover a variety of astrodynamics applications, including the unperturbed and perturbed two-body problem, as well as the mapping of statistical moments within the framework of the circular restricted three-body problem. Our results show how our method consistently propagates non-Gaussian uncertainties with contained computational cost while maintaining low errors in comparison to statistical moments generated via more traditional Monte Carlo simulations.
Notably, the high-order moment computation is performed via a symbolic manipulator independently of the dynamics and for various probability density functions, effectively reducing the computational burden of uncertainty propagation to only the calculation of Taylor series coefficients around a nominal trajectory. This is performed by efficiently integrating the variational equations of the system.
Future work will focus on exploring the opportunities enabled by our tool for trajectory design and stochastic optimization under uncertain initial conditions and mission parameters.

### v2

@@ 3.1.2 Perturbed Case
$\dot{\boldsymbol{r}}$ $=\boldsymbol{v}$
$\dot{v}_{x}$ $=-\frac{\mu}{r^{3}}x-\dfrac{3\mu J_{2}R_{e}^{2}}{2r^{5}}\bigg{(}1-\dfrac{5z^{2}}{r^{2}}\bigg{)}x$
$\dot{v}_{y}$ $=-\frac{\mu}{r^{3}}y-\dfrac{3\mu J_{2}R_{e}^{2}}{2r^{5}}\bigg{(}1-\dfrac{5z^{2}}{r^{2}}\bigg{)}y$
$\dot{v}_{z}$ $=-\frac{\mu}{r^{3}}z-\dfrac{3\mu J_{2}R_{e}^{2}}{2r^{5}}\bigg{(}3-\dfrac{5z^{2}}{r^{2}}\bigg{)}z\text{,}$
@@
As for the unperturbed case, also in this case the coupled equations of motion, cause the initial uncertainties in the $x,y$ component and in the parameters to eventually affect all six components of the state during the propagation. We use the Monte Carlo propagation with 10 million samples as ground truth, and we compare the reconstructed statistical moments with those found with the proposed approach, up to the Taylor expansion of order 5. As before, the time required to compute the Taylor expansion at order five of the state is orders of magnitude lower than the cost to perform the Monte Carlo. On the left of Fig. 4, we display how the accuracy of the reconstruction in the first, second, and third moment behaves as a function of time, using the absolute and relative errors in the moment reconstruction discussed in Eq. (24). The error is always maintained below 1% of accuracy, for all the first three moments tensors. Furthermore, as seen from the right figure, order 1, which consists of just using the first-order state transition matrix, does not accurately reconstruct the second moment. Instead, increasing the order to two, three, and four, brings substantial improvements in the error of the reconstruction, while between order four and order five there is not a significant gap.
Figure 4: Errors in the reconstruction of the first three moments (left), and of the second moment for different orders, both as a function of time (right).
@@ 3.2.1 Taylor Polynomial Convergence Radius of Different Families
Table 2: Initial conditions, period, mass ratio parameter, and stability for the CR3BP systems and periodic orbits used.
@@
Figure 5: Three-dimensional CR3BP periodic orbits in the Earth-Moon (upper left) and Enceladus-Saturn (upper center) systems and Cauchy-Hadamard’s radius of convergence estimation.
To generate these, for each orbit, we construct the Taylor expansion after a period and compute the Cauchy-Hadamard radius of convergence for each component of the state. Then, the lowest radius across the different components estimated at the highest order can be used as a proxy of the ball radius in which the Taylor series converges. Hence, this guides the choice for the initial uncertainty size that can be acceptable. While this criterion is often used in practice, it is important to remember that it would only be valid if the Taylor series is expanded to infinite terms: truncating at lower orders only results in an approximation of the convergence radius. To verify that the criterion is working properly, in Fig. 6, we also report two projections of the ground truth Poincaré map, propagated at different ball radii, together with the associated error w.r.t. the Taylor approximation, for all the different periodic orbits (each row corresponds to a periodic orbit). In the figure, we use the same color for ball radii of the same size. As we observe, the error of the Taylor approximation is maintained very low, when the perturbation is below the Cauchy-Hadamard’s radius of convergence. We also perform the exact same analysis for the Saturn-Enceladus short-period L4 orbit, which we report in Fig. 7. Also in this case, we confirm that for perturbations below the Cauchy-Hadamard’s radius of convergence, the errors between the Taylor approximation and the ground truth orbit are maintained very low, thereby confirming that this criterion can be used as a good proxy for the region of validity of the Taylor expansion.
Figure 6: 1st and 3rd column: numerically integrated Poincaré maps; 2nd and 4th column: error between the ground truth and the Taylor approximation.
@@ 3.2.2 Statistical Moments Mapping onto Poincaré Maps
We first compute the relative error between the reconstructed and ground truth second order moment using Eq. (24), obtaining for order 1 $\varepsilon_{r,2}=2.4017$, for order 2 $\varepsilon_{r,2}=0.0109$, for order 3 $\varepsilon_{r,2}=0.0106$, for order 4, and order 5, $\varepsilon_{r,2}=0.0105$. This confirms not only that the accuracy improves, as higher orders for the Taylor expansion are used, but also that for this case there is not an improvement between order 4 and order 5, which suggests that order 4 might have sufficed if the objective was the reconstruction of the second moment only.
@@ 4 Conclusion
In this Note, we present a novel approach for uncertainty propagation using high-order Taylor expansions of the flow and moment generating functions. While previous studies predominantly focus on Gaussian distributions, our method leverages the relationship between moment generating functions and distribution moments to extend high-order uncertainty propagation techniques to non-Gaussian scenarios, broadening their applicability across diverse problems and uncertainty types.
The proposed approach is general by design and suitable for arbitrary dynamical systems, as demonstrated through numerical simulations across various astrodynamics applications, including the unperturbed and perturbed two-body problem and the mapping of statistical moments within the framework of the circular restricted three-body problem. Our results reveal that the method effectively propagates non-Gaussian uncertainties with manageable computational costs and achieves low error rates compared to statistical moments generated via traditional Monte Carlo simulations.
Notably, high-order moment computation is performed once, using a symbolic manipulator independent of specific dynamics and probability density functions. This optimization reduces the computational burden of uncertainty propagation to the calculation of Taylor series coefficients around a nominal trajectory, which is efficiently executed through the integration of the system’s variational equations. The code supporting this work is available open-source [^2] https://gitlab.surrey.ac.uk/stag/up-with-mgf, accessed July 2024
, to facilitate researchers to use and build upon these techniques.